OpenStudy (anonymous):

The measures of the angles in a quadrilateral are represented by x, 2x, 3x, and 3x. Write an equation that would allow you to solve for the value of x. Solve for the value of x

OpenStudy (welshfella):

the 4 angles in a quadrilateral add up to 360 degrees so we have x +_ 2x + 3x + 3x = 360

OpenStudy (mathstudent55):

We need the same formula again. Use the formual with n = 4. What do you get for S?

OpenStudy (mathstudent55):

\(S = (n - 2)180\) In this problem, n = 4 since a quadrilateral has 4 sides.

OpenStudy (anonymous):

I got 360

OpenStudy (mathstudent55):

Good. The sum of the measures of the angles is 360.

OpenStudy (mathstudent55):

We are told the 4 angles measure: x, 2x, 3x, and 3x. If we add up those measures, we get: \(x + 2x + 3x + 3x\) We know the angle measures must add up to 360 from the formula, so that means \(x + 2x + 3x + 3x = 360\) Now we can solve this equation for x.

OpenStudy (mathstudent55):

We add all the terms on the left side: \(9x = 360\) Now we divide both sides by 9: \(x = 40\)